OpenStudy (anonymous):

If a% of 180 is 95. Find a.

hartnn (hartnn):

% means divide by 100 so, \(\large \frac{a} {100}\times 180 = 95\) now can u find a=... ?

OpenStudy (anonymous):

is the answer equals to 475/9?

hartnn (hartnn):

yes! exactly... good :)

OpenStudy (anonymous):

thank you so much for your help!

hartnn (hartnn):

welcome ^_^

hartnn (hartnn):

bdw, \(\huge \color{red}{\text{Welcome to Open Study}}\ddot\smile\)

OpenStudy (anonymous):

It is very easy a/100x180=95 180a=95x100 180a=9500 a=9500/180 a=52.77778